A six faced dice has two red faced, three blue and one green face. It is thrown twice. What is the chance that red face appear in both the throws?
Answer by sudhanshu_kmr(1152)


probability that red face appear on single throw = 2/6

probability that red face appear on both throws = 2/6 * 2/6
= 4/36
= 1/9